# AutoCodeRover: Autonomous Program Improvement

### v1

@@ Abstract.
Researchers have made significant progress in automating the software development process in the past decades. Automated techniques for issue summarization, bug reproduction, fault localization, and program repair have been built to ease the workload of developers. Recent progress in Large Language Models (LLMs) has significantly impacted the development process, where developers can use LLM-based programming assistants to achieve automated coding. Nevertheless software engineering involves the process of program improvement apart from coding, specifically to enable software maintenance (e.g. program repair to fix bugs) and software evolution (e.g. feature additions). In this paper, we propose an automated approach for solving Github issues to autonomously achieve program improvement (program repair and feature addition). In our approach called AutoCodeRover, LLMs are combined with sophisticated code search capabilities, ultimately leading to a program modification or patch. In contrast to recent LLM agent approaches from AI researchers and practitioners, our outlook is more software engineering oriented. We work on a program representation (abstract syntax tree) as opposed to viewing a software project as a mere collection of files. Our code search exploits the program structure in the form of classes/methods to enhance LLM’s understanding of the issue’s root cause, and effectively retrieve a context via iterative search. The use of spectrum based fault localization using tests, further sharpens the context, as long as a test-suite is available. We consider the recently proposed SWE-bench-lite which consists of 300 real-life Github issues involving bug fixing and feature additions. Experiments on SWE-bench-lite show increased efficacy in solving Github issues (more than 20% on SWE-bench-lite), as compared to recent efforts from the AI community. Interestingly, our approach resolved 67 GitHub issues in less than ten minutes each, whereas developers spent more than 2.77 days on average. We posit that our workflow enables autonomous software engineering, where, in future, auto-generated code from LLMs can be autonomously improved.
@@ 1. Beyond Automatic Programming
Automating software engineering tasks has long been a vision among software engineering researchers and practitioners. One of the key challenges has been the handling of ambiguous natural language requirements, in the process of automatic programming. In addition, there has been progress in some other software engineering activities such as automated test generation via fuzzing / symbolic execution (Böhme et al., 2021; Cadar and Sen, 2013), automated program repair (Goues et al., 2019) via search / symbolic execution, and so on.
Recent progress in large language models (LLMs) and the appearance of tools like Github Copilot (Ovalle, 2023) hold significant promise in automatic programming. These tools show the possibility of converting natural language requirements into full-blown programs for a variety of examples. This progress immediately raises the question of whether such automatically generated code can be trusted to be integrated into software projects, and if not, what improvements to the technology are needed. One possibility is to automatically repair generated code to achieve trust. This brings out the importance of automating program repair tasks towards achieving the vision of autonomous software engineering.
Given this motivation of automating program repair, and the large number of hours developers often spend manually fixing bugs, we looked into the possibility of fully autonomous program repair. Specifically, we feel that bug fixing and feature addition are the two key categories of tasks that a development team may focus on when maintaining an existing software project. To achieve this goal, we proposed an approach that augments LLM with context knowledge from the code repository. We call our tool AutoCodeRover since it iteratively roams across the codebase and enriches LLM towards a better understanding of particular GitHub issues by exploiting the program structure.
@@
AutoCodeRover also can leverage popular debugging techniques such as spectrum-based fault localization (SBFL) (Wong et al., 2016a) to decide more precise code search APIs for context retrieval if a test suite accompanying the project is available. SBFL primarily considers the control flow of the passing and failing tests and assigns a suspiciousness score to the different methods of the program. The LLM agent may prioritize retrieving context from particular methods and classes if the fault localization result is provided, e.g., when a method appears both in the issue description and in the output of fault localization. In the last step, AutoCodeRover may perform patch validation (if test-suite available) to determine whether the patch produced by AutoCodeRover passes the tests in the given test-suite. Otherwise, AutoCodeRover will rerun the patch generation with a retry threshold until a correct patch is found.
We see the need to endow a software engineering oriented outlook to the recent flurry of activity on LLM agents for software engineering, which mostly has only an AI flavor. We believe that the software engineering angle can be conveyed via four dimensions, and can enhance the capabilities of LLM agents.
  • We work on program representations (abstract syntax tree or AST) as opposed to viewing a software project as a collection of files. We posit that working on program representations like AST will be critical to autonomous software engineering workflows.
@@
  • We posit that higher efficacy of automated repair is more important than time efficiency, as long as the time is within a threshold. It is well-known in empirical software engineering research that time limits of 30 -60 minutes for automated repair are tolerable, based on extensive developer surveys in the field (Noller et al., 2022). We thus set a modest timelimit of 10 minutes and report 22% efficacy on SWE-bench-lite in solving Gihub issues. We can compare this timelimit of 10 minutes to the average time of 2.77 days to fix the Github issues manually.
  • Finally, one should be able to exploit debugging/ analysis techniques like test-based fault localization to guide the search for code in resolving Github issues. We show that use of fault localization in setting the code context leads to an increase in efficacy from 16% - 20% in a single run of AutoCodeRover on SWE-bench-lite.
Our contribution lies in the effective use of code search to make software engineering processes like program repair autonomous. In a forward-looking perspective, we show that our approach can be used to automate similar program improvement tasks - such as feature addition. We report favorable experimental results (22% efficacy in resolving Github issues) on SWE-bench-lite. AutoCodeRover helps in making software engineering autonomous - and it can work on both manually written code as well as automatically generated code. This can help lead us to future day programming environments – which may involve autonomous improvement of automatically generated code.
2. Background
@@ 2.1. Program Repair
Test-suite based automated program repair (APR) has attracted significant attention in the last decade (Goues et al., 2019). These techniques aim to generate a patch for a buggy program to pass a given test-suite. APR techniques typically include search-based, semantic-based, and pattern/learning-based APR. Search-based APR techniques like GenProg (Weimer et al., 2009) take a buggy program and generate patches using predefined code mutation operators, or search for a patch over the patch space that passes the given test suite. Semantics-based APR techniques (Nguyen et al., 2013; Mechtaev et al., 2016) generate patches by formulating a repair constraint that needs to be satisfied based on a given test suite specification, and then solving the repair constraint to generate patches. Learning-based APR techniques (Lutellier et al., 2020; Smith et al., 2015; Zhu et al., 2021) often train a deep learning model with large code repositories and are guided by a specific representation of code syntax and semantics to predict the next tokens that are most likely to be correct patch.
Recent work (Fan et al., 2023; Jiang et al., 2023; Xia et al., 2023) have shown use of LLMs for automated program repair. This line of work often assumes the buggy program statements are given (i.e. perfect fault localization assumption) and focuses on constructing APR-specific prompts that guide LLM to generate a patch for the selected buggy program statements multiple times until a patch that passes all tests is found. However, obtaining buggy locations for a large project is an essential and challenging task in resolving real-life bug reports.
@@ 2.2. Dataset
SWE-bench (Jimenez et al., 2024) is a benchmark that aims to evaluate the capabilities of large language models in resolving end-to-end real-life software engineering tasks. The benchmark consists of 2,294 real-life software engineering task instances collected from the repositories of 12 popular large Python projects  [^1] https://www.swebench.com/index.html (e.g., django, sympy). Each SWE-bench task instance contains a pair of Github issue and corresponding pull requests. The GitHub issue either reports a bug to be fixed or requests to implement new features. The pull request includes the code changes made by human developers to resolve the issue and test cases that prevent the issue. Unlike traditional code generation tasks in HumanEval (Chen et al., 2021) and MBPP (Li et al., 2022) benchmark, resolving a SWE-bench instance is particularly challenging because it requires automatically generating code changes that address the problem in a GitHub issue for a matured large code repository based only on the issue description. More specifically, the process may involve a series of complex tasks like reasoning the target bug location across files in the code repository, analyzing the root cause of the issue, proposing bug-fixing strategies, and eventually writing a patch that passes all the test cases added in the pull request.
SWE-bench team also constructed a lite version of SWE-bench (SWE-bench lite) due to the high cost of evaluating in the complete SWE-bench. SWE-bench lite includes 300 task instances sampled from SWE-bench, following a similar Github issue repository distribution as SWE-bench. SWE-bench lite removed task instances with meta files such as images, external links, and instances without proper description (e.g., short problem statement fewer than 40 words). In the following sections, we describe our design of an AI program improvement framework for handling real-life software issues and demonstrate its effectiveness in SWE-bench lite.
@@ 3. Motivating Example
In this section, we illustrate how our tool AutoCodeRover can collect code context and generate a patch from an issue description and the corresponding project codebase. We show an example of a challenging feature addition task. Figure 1 demonstrates the workflow of AutoCodeRover on an issue submitted to the Django issue tracker [^2] https://code.djangoproject.com/ticket/32347. This issue is classified as “New feature” in the issue tracker, and is included in SWE-bench lite with the id “django-13933”. The first part of Figure 1 shows the issue description (the code part is simplified for brevity). This issue requests adding support to the ModelChoiceField class, so that it “shows the value of the invalid choice when raising a validation error”.
@@
In Step 3, the agent receives the implementation of validate and to_python methods. Among the two methods, it selects to_python to be the more suitable place to make changes, since to_python raises the relevent exception and does not include the invalid value. At this point, the retrieval agent deems the collected code context as sufficient for understanding the issue and drafting the patch. The identified buggy location, together with the gathered context and analysis so far, is passed to another patch generation agent. This agent is instructed to write patches following a specified format. In Step 4, a patch is written to allow a value to be integrated into the error message, by utilizing %-formatting in Python. Although written in a different way compared to the developer patch, this patch achieves similar effect and passes the developer provided test-suite for this issue.
@@ 4. AI Program Improvement Framework
In this section, we discuss the design of AutoCodeRover. AutoCodeRover is a system incorporating AI agents for program improvement tasks in large software projects. AutoCodeRover is designed to work in a realistic software development lifecycle, in which users submit issue reports to a software repository describing a bug, and the project maintainers craft a patch to resolve the issue. With a submitted issue, AutoCodeRover autonomously analyzes the submitted issue, retrieves the relevant code context in the software project, and generating a patch. This patch can then be verified by human developers, thereby reducing manual efforts.
@@
search_class (cls)
Search for class cls in the codebase.
Signature of the searched class.
search_class_in_file (cls, f)
Search for class cls in file f.
Signature of the searched class.
search_method (m)
Search for a method in the codebase.
Implementation of the searched method.
search_method_in_class (m, cls)
Search for method m in class cls.
Implementation of the searched method.
search_method_in_file (m, f)
Search for a method m in file f.
Implementation of the searched method.
search_code (c)
Search for code snippet c in the codebase.
Region of code surrounding the searched snippet.
search_code_in_file (c, f)
Search for code snippet c in file f.
Region of code surrounding the searched snippet.
@@ 4.1. Overview
We first describe the overall stages AutoCodeRover operates in, and will proceed in more details in the rest of this section.
The overall workflow of AutoCodeRover is shown in Figure 2. AutoCodeRover takes in as input a problem statement $P$ of the issue to be resolved, and a codebase $C$ of the corresponding software project. This problem statement $P$ contains the title and description of the issue, as shown in Section 3. From a problem statement written in natural language, AutoCodeRover analyzes the requirement from it and proceeds in two main stages, which are context retrieval and patch generation.
In the context retrieval stage, AutoCodeRover employs an LLM agent to navigate through a potentially large codebase $C$ and extract the relevant code snippets relevant to $P$. This navigation is enabled by a set of context retrieval APIs (Section 4.2), which enables an LLM to retrieve information about the project (e.g. class signatures) and actual code snippets (e.g. method implementations). The context retrieval agent directs this navigation, and decides which retrieval APIs to use based on the current available context. In order to make the LLM-directed navigation more “controlled”, we devise a stratified stategy of invoking the retrieval APIs (Section 4.3). This stratified strategy instructs the LLM to only invoke necessary retrieval APIs based on the available information, and iteratively changes the set of retrieval APIs used when more code-related context are returned from the previous API calls.
@@
We note that the workflow of AutoCodeRover discussed so far only requires the problem statement $P$ and codebase $C$ as input, and do not require any program specifications such as testcases. However, when testcases are available (e.g. provided by developers or generated from another system), execution-based information and program analysis techniques can be integrated into the AutoCodeRover framework (Section 4.4). For example, a statistical fault localization (Wong et al., 2016a) tool can be used to reveal more relevant methods beyond those mentioned in the problem statement. The additionally revealed code context can thereby influence the set of retrieval APIs invoked in the context retrieval stage. Moreover, with testcases available, the patch generation agent can employ an additional patch validation step in its retry-loop when crafting a patch. In the remainder of this section, we discuss in greater detail the components of AutoCodeRover and their design considerations.
@@ 4.2. Context Retrieval APIs
In the context retrieval stage, the basic components are a set of APIs which an LLM agent can use to gather relevant code context and snippets from the codebase. For typical software engineering issues, we observe that users often metion some “hints” on which part of the codebase is relevant. These hints can be the names of the relevant methods, classes, or files, and sometimes also contain short code snippets. Although these hints may not directly point to the precise location for code modification, they often reveal code context in the project that is relevant to the current issue. Figure 4 illustrates two real-world issues submitted to the sympy and django projects, respectively. The “hints” contained in these two issue descriptions are highlighted. In sympy-12481, the class Permutation is mentioned twice; in django-13230, multiple hints are mentioned, such as code snippet feed.add_item(), package path syndication.views, method name item_extra_kwargs and etc.
Based on the type of project and code related hints, we design a set of APIs for an LLM agent to retrieve code context from these hints. [^3] Forward calculation of program dependencies from the hints (Jin et al., 2024) can be potentially used at this stage, to further aid the retrieval. The current set of APIs in AutoCodeRover and their outputs are shown in Table 1. Once invoked by the LLM agent, the retrieval APIs search for classes, methods and code snippets in the codebase, and return the researched results back to the agent. The outputs from all the API invocations form the relevant code context for patch generation later on. To avoid forming very lengthy code context that creates distraction during patch generation, we return only necessary information as API outputs. For example, since the complete definition of a class can be lengthy in large projects, we only return signature of the class as output for search_class and search_class_in_file. This kind of returning of signature is a better approach than cutting off the context at a certain bound. Upon receiving the class signature, the agent could then invoke another API to search for the relevant methods / snippets inside the class.
@@
For the LLM agent to invoke the context retrieval APIs, the description and expected output of them are presented to it as part of prompt. When the agent decides to invoke a set of retrieval APIs, it responds with the list of API call names and the corresponding arguments. These retrieval API requests are processed locally by parsing a local codebase of the project into Abstract Syntax Tree (AST) and searching over it. Results of locally executing these APIs are returned to the agent, forming the code context.
Comparing to other retrieval methods.
An alternative for gathering relevant code context is to utilize information retrieval based methods such as BM25 retrieval to find the relevant files in the project. The BM25 algorithm (ROBERTSON et al., 1995; Robertson et al., 2009) ranks the set of files in the project based on their relevance to the words appeared in the issue description. In software engineering tasks, general retrieval methods such as BM25 retrieval can be less effective. Many content in the issue decription, such as text describing the reproduction scenario, may not be directly relevant to the code in the project. Instead, our method utilizes LLM to identify code-related keywords in the natural languange descriptions, and performs code-specific search by traversing the AST.

### v2

@@ Abstract.
Researchers have made significant progress in automating the software development process in the past decades. Automated techniques for issue summarization, bug reproduction, fault localization, and program repair have been built to ease the workload of developers. Recent progress in Large Language Models (LLMs) has significantly impacted the development process, where developers can use LLM-based programming assistants to achieve automated coding. Nevertheless software engineering involves the process of program improvement apart from coding, specifically to enable software maintenance (e.g. program repair to fix bugs) and software evolution (e.g. feature additions). In this paper, we propose an automated approach for solving Github issues to autonomously achieve program improvement. In our approach called AutoCodeRover, LLMs are combined with sophisticated code search capabilities, ultimately leading to a program modification or patch. In contrast to recent LLM agent approaches from AI researchers and practitioners, our outlook is more software engineering oriented. We work on a program representation (abstract syntax tree) as opposed to viewing a software project as a mere collection of files. Our code search exploits the program structure in the form of classes/methods to enhance LLM’s understanding of the issue’s root cause, and effectively retrieve a context via iterative search. The use of spectrum based fault localization using tests, further sharpens the context, as long as a test-suite is available. We consider the recently proposed SWE-bench-lite which consists of 300 real-life Github issues involving bug fixing and feature additions. Experiments on SWE-bench-lite show increased efficacy in solving Github issues (22-23% on SWE-bench-lite), as compared to recent efforts from the AI community. Interestingly, our approach resolved 67 GitHub issues in less than 12 minutes each, whereas developers spent more than 2.77 days on average. On the full SWE-bench consisting of 2294 GitHub issues, AutoCodeRover solved around 16% of issues, which is higher than the efficacy of the recently reported AI software engineer Devin from Cognition Labs, while taking time comparable to Devin. We posit that our workflow enables autonomous software engineering, where, in future, auto-generated code from LLMs can be autonomously improved.
@@ 1. Beyond Automatic Programming
Automating software engineering tasks has long been a vision among software engineering researchers and practitioners. One of the key challenges has been the handling of ambiguous natural language requirements, in the process of automatic programming. In addition, there has been progress in some other software engineering activities such as automated test generation (Böhme et al., 2021; Cadar and Sen, 2013), automated program repair (Goues et al., 2019), and so on.
Recent progress in large language models (LLMs) and the appearance of tools like Github Copilot (Ovalle, 2023) hold significant promise in automatic programming. This progress immediately raises the question of whether such automatically generated code can be trusted to be integrated into software projects, and if not, what improvements to the technology are needed. One possibility is to automatically repair generated code to achieve trust. This brings out the importance of automating program repair tasks towards achieving the vision of autonomous software engineering.
Given this motivation of automating program repair, and the large number of hours developers often spend manually fixing bugs, we looked into the possibility of fully autonomous program improvement. Specifically, we feel that bug fixing and feature addition are the two key categories of tasks that a development team may focus on when maintaining an existing software project. To achieve this goal, we proposed an approach that augments LLM with context knowledge from the code repository. We call our tool AutoCodeRover.
@@
AutoCodeRover also can leverage debugging techniques such as spectrum-based fault localization (SBFL) (Wong et al., 2016a) to decide more precise code search APIs for context retrieval if a test suite accompanying the project is available. SBFL primarily considers the control flow of the passing and failing tests and assigns a suspiciousness score to the different methods of the program. The LLM agent may prioritize retrieving context from particular methods and classes if the fault localization result is provided, e.g., when a method appears both in the issue description and in the output of fault localization. In the last step, AutoCodeRover may perform patch validation using available tests, to determine whether the patch produced by AutoCodeRover passes the tests in the given test-suite. Otherwise, AutoCodeRover will rerun the patch generation with a retry limit until a correct patch is found.
Contributions
Our contribution lies in the effective use of code search to make software engineering processes like program repair autonomous. We demonstrate this capability by autonomously solving GitHub issues. We report favorable experimental results on the SWE-bench dataset (Jimenez et al., 2024b). We achieve 15.95% efficacy on the full SWE-bench with 2294 GitHub issues, and 22.33% efficacy on the SWE-bench-lite subset with 300 GitHub issues. Overall, we see the need to endow a software engineering oriented outlook to the recent flurry of activity on LLM agents for software engineering, which mostly has only an AI flavor. We believe that the software engineering angle can be conveyed via (at least) the following five dimensions, and can enhance the capabilities of LLM agents.
  • We work on program representations (abstract syntax tree or AST) as opposed to viewing a software project as a collection of files. We posit that working on program representations like AST will be useful in autonomous software engineering workflows.
@@
  • We posit that higher efficacy of automated repair is more important than time efficiency, as long as the time is within a threshold. It is well-known in empirical software engineering research that time limits of 30-60 minutes for automated repair are tolerable, based on extensive developer surveys in the field (Noller et al., 2022). We thus report 22% efficacy on SWE-bench-lite in solving GitHub issues, within 12 minutes. We can compare this timelimit of 12 minutes to the average time of 2.77 days to fix the GitHub issues manually.
  • One should be able to exploit debugging techniques like test-based fault localization to guide the search for code in resolving GitHub issues. We show that use of fault localization in setting the code context leads to an increase in efficacy of AutoCodeRover in solving GitHub issues.
  • Finally, it is also useful to examine how many of the solved GitHub issues are producing acceptable patches. We study the patches produced by AutoCodeRover and report that 2/3 of the autonomously produced patches from AutoCodeRover are correct and acceptable. We note that this aspect has not been reported by Devin (Labs, 2024) and SWE-agent (Yang et al., 2024).
2. Relevant Literature
@@ 2.1. Program Repair
Test-suite based automated program repair (APR) has attracted significant attention in the last decade (Goues et al., 2019). These techniques aim to generate a patch for a buggy program to pass a given test-suite. APR techniques typically include search-based, semantic-based, and pattern/learning-based APR. Search-based APR techniques like GenProg (Weimer et al., 2009) take a buggy program and generate patches using predefined code mutation operators, or search for a patch over the patch space that passes the given test suite. Semantics-based APR techniques (Nguyen et al., 2013; Mechtaev et al., 2016) generate patches by formulating a repair constraint that needs to be satisfied based on a given test suite specification, and then solving the repair constraint to generate patches. Learning-based APR techniques (Lutellier et al., 2020; Smith et al., 2015; Zhu et al., 2021) often train a deep learning model with large code repositories and are guided by a specific representation of code syntax and semantics to predict the next tokens that are most likely to be correct patch. There are also works (Koyuncu et al., 2019; Tan et al., 2024; Motwani and Brun, 2023) that tried to leverage GitHub issues and bug reports to improve APR effectiveness.
Recent work (Fan et al., 2023; Jiang et al., 2023; Xia et al., 2023) have shown the use of LLMs for automated program repair. This line of work often assumes the buggy program statements are given (i.e. perfect fault localization assumption) and focuses on constructing APR-specific prompts that guide LLM to generate a patch for the selected buggy program statements multiple times until a patch that passes all tests is found. However, obtaining buggy locations for a large project is an essential and challenging task in resolving real-life bug reports.
@@ 2.2. Dataset
SWE-bench (Jimenez et al., 2024b) is a benchmark that aims to evaluate the capabilities of large language models in resolving end-to-end real-life software engineering tasks. The benchmark consists of 2,294 real-life software engineering task instances collected from the repositories of 12 popular large Python projects  [^1] https://www.swebench.com/index.html (e.g., django, sympy). Each SWE-bench task instance contains a pair of Github issue and corresponding pull requests. The GitHub issue either reports a bug to be fixed or requests to implement new features. The pull request includes the code changes made by human developers to resolve the issue and test cases that prevent the issue. Unlike traditional code generation tasks in HumanEval (Chen et al., 2021) and MBPP (Li et al., 2022) benchmark, resolving a SWE-bench instance is particularly challenging because it requires automatically generating code changes that address the problem in a GitHub issue for a matured large code repository based only on the issue description. More specifically, the process may involve a series of complex tasks like reasoning the target bug location across files in the code repository, analyzing the root cause of the issue, proposing bug-fixing strategies, and eventually writing a patch that passes all the test cases added in the pull request.
SWE-bench team also constructed a lite version of SWE-bench (SWE-bench lite) due to the high cost of evaluating in the complete SWE-bench. SWE-bench lite includes 300 task instances sampled from SWE-bench, following a similar repository distribution. SWE-bench lite removed task instances with meta files such as images, external links, and instances without proper description (e.g., short problem statement which is fewer than 40 words).
@@ 3. Motivating Example
In this section, we illustrate how our tool AutoCodeRover can collect code context and generate a patch from an issue description and the corresponding project code-base. We show an example of a feature addition task. Figure 1 demonstrates the workflow of AutoCodeRover on an issue submitted to the Django issue tracker [^2] https://code.djangoproject.com/ticket/32347. This issue is classified as “New feature” in the issue tracker, and is included in SWE-bench lite with the id “django-13933”. The first part of Figure 1 shows the issue description (the code part is simplified for brevity). This issue requests adding support to the ModelChoiceField class, so that it “shows the value of the invalid choice when raising a validation error”.
@@
In Step 3, the agent receives the implementation of validate and to_python methods. Among the two methods, it selects to_python to be the more suitable place to make changes, since to_python raises the relevent exception and does not include the invalid value. At this point, the retrieval agent deems the collected code context as sufficient for understanding the issue and drafting the patch. The identified buggy location, together with the gathered context and analysis so far, is passed to another patch generation agent. This agent is instructed to write patches following the format described in Step 4 (see yellow box) of Figure 2. In Step 4, a patch is written to allow a value to be integrated into the error message, by utilizing %-formatting in Python. Although written in a different way compared to the developer patch, this patch achieves similar effect and passes the developer provided test-suite for this issue.
@@ 4. AI Program Improvement Framework
In this section, we discuss the design of AutoCodeRover. AutoCodeRover is a system incorporating AI agents for program improvement tasks in large software projects. AutoCodeRover is designed to work in a realistic software development lifecycle, in which users submit issue reports to a software repository describing a bug, and the project maintainers craft a patch to resolve the issue. With a submitted issue, AutoCodeRover autonomously analyzes the submitted issue, retrieves the relevant code context in the software project, and generates a patch. This patch can then be vetted by human developers. If such a tool can automatically handle a certain percentage of the issues awaiting developers’ attention, manual efforts are reduced.
@@
search_class (cls) Search for class cls in the codebase. Signature of the searched class.
search_class_in_file (cls, f) Search for class cls in file f. Signature of the searched class.
search_method (m) Search for method m in the codebase. Implementation of the searched method.
search_method_in_class (m, cls) Search for method m in class cls. Implementation of the searched method.
search_method_in_file (m, f) Search for method m in file f. Implementation of the searched method.
search_code (c) Search for code snippet c in the codebase. Region of code surrounding the searched snippet.
search_code_in_file (c, f) Search for code snippet c in file f. Region of code surrounding the searched snippet.
@@ 4.1. Overview
We first describe the overall stages AutoCodeRover operates in, and will proceed in more details in the rest of this section. The overall workflow of AutoCodeRover is shown in Figure 2. AutoCodeRover takes in as input a problem statement $P$ of the issue to be resolved, and a codebase $C$ of the corresponding software project. This problem statement $P$ contains the title and description of the issue, as shown in Section 3. From a problem statement written in natural language, AutoCodeRover analyzes the requirement from it and proceeds in two main stages, which are context retrieval and patch generation.
In the context retrieval stage, AutoCodeRover employs an LLM agent to navigate through a potentially large codebase $C$ and extract the relevant code snippets relevant to $P$. This navigation is facilitated by a set of context retrieval APIs (Section 4.2), which enables an LLM to retrieve information about the project (e.g. class signatures) and actual code snippets (e.g. method implementations). The context retrieval agent directs this navigation, and decides which retrieval APIs to use based on the current available context. In order to make the LLM-directed navigation more “controlled”, we devise a stratified stategy of invoking the retrieval APIs (Section 4.3). This stratified strategy instructs the LLM to only invoke necessary retrieval APIs based on the available information, and iteratively changes the set of retrieval APIs used when more code-related context are returned from the previous API calls.
@@
We note that the workflow of AutoCodeRover discussed so far only requires the problem statement $P$ and codebase $C$ as input, and do not require any program specifications such as testcases. However, when testcases are available (e.g. provided by developers or generated from another tool), execution-based information and program analysis techniques can be integrated into the AutoCodeRover framework (Section 4.4). For example, statistical fault localization (Wong et al., 2016a) tool can be used to reveal more relevant methods beyond those mentioned in the problem statement. The additionally revealed code context can thereby influence the set of retrieval APIs invoked in the context retrieval stage. Moreover, with testcases available, the patch generation agent can employ an additional patch validation step in its retry-loop when crafting a patch. In the remainder of this section, we discuss in greater detail the components of AutoCodeRover and their design considerations.
@@ 4.2. Context Retrieval APIs
In the context retrieval stage, the basic components are a set of APIs which an LLM agent can use to gather relevant code context and snippets from the codebase. For typical software project issues, we observe that users often mention some “hints” on which part of the codebase is relevant. These hints can be the names of the relevant methods, classes, or files, and sometimes also contain short code snippets. Although these hints may not directly point to the precise location for code modification, they often reveal code context in the project that is relevant to the current issue. Figure 4 illustrates two real-world issues submitted to the sympy and django projects, respectively, and the “hints” are highlighted. In sympy-12481, the class Permutation is mentioned twice; in django-13230, multiple hints are mentioned, such as code snippet feed.add_item(), package path syndication.views, method name item_extra_kwargs and etc. Based on the type of project and code related hints, we design a set of APIs for an LLM agent to retrieve code context from these hints. The current set of APIs in AutoCodeRover and their outputs are shown in Table 1. Once invoked by the LLM agent, the retrieval APIs search for classes, methods and code snippets in the code-base, and return the results back to the agent. To avoid forming very lengthy code context that creates distraction during patch generation, we return only necessary information as API outputs. For example, since the complete definition of a class can be lengthy in large projects, we only return signature of the class as output for search_class and search_class_in_file. Returning the signature to shorten cntext, is a better approach than cutting off the context at a certain bound. Upon receiving the class signature, the agent could then invoke another API to search for the relevant methods / snippets inside the class.
@@
For the LLM agent to invoke the context retrieval APIs, the description and expected output of them are presented to it as part of prompt. When the agent decides to invoke a set of retrieval APIs, it responds with the list of API call names and the corresponding arguments. These retrieval API requests are processed locally by parsing a local codebase of the project into AST and searching over it. Results of locally executing these APIs are returned to the agent, forming the code context.